Circular No. 8031

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jo IN NGC 5708 Further to IAUC 7373, A. Dimai reports the discovery of an apparent

supernova (mag 16.4) on unfiltered CCD images taken on Dec. 11.20 and 11.23 UT, confirmed at mag 16.3 on an image taken on Dec. 12.21. The new object is located at α = 14^{h}38^{m}16^{s}.42, δ = +40^{o}27′14′′.0 (equinox 2000.0), which is 1′′ east and 10′′ south of the nucleus of NGC 5708. SN 2002jo is not present in Dimai’s CCD images from Sept. 5 (limiting mag 18.5) or on Palomar Sky Survey red and blue digitized plates.

R. J. Foley, M. Papenkova, R. Chornock, W. D. Li, S. Jha, and A. V. Filippenko, University of California, Berkeley, report that spectra (range 310–1000 nm), obtained on Dec. 12 UT with the Shane 3-m telescope at Lick Observatory, show this to be a supernova of type Ia, ∼ 3 weeks past maximum light. The continuum is redder than that of SN 1994D at a similar epoch, possibly indicating that the object is significantly extinguished by dust.

SUPERNOVA 2002jp IN NGC 3313 Further to IAUC 8030, M. Ganeshalingam and W. Li report the dis-

covery by LOTOSS of a supernova on unfiltered images, with the following photometry available (via KAIT, except Dec. 10.5 UT by M. Schwartz with the Tenagra II telescope): May 14.2, [18.5; Nov. 23.5, 16.9; Dec. 7.5, 16.7; 10.5, 17.1:; 11.5, 17.2. SN 2002jp is located at α = 10^{h}37^{m}26^{s}.50, δ = −25^{o}21′35′′.2 (equinox 2000.0), which is 13′′.6 east and 145′′.9 south of NGC 3313. Foley et al. report that spectra, taken as above, show that SN 2002jp is a type-Ic supernova, ∼ 2 weeks past maximum light.

SUPERNOVAE 2002jj AND 2002jm Foley et al. also write that inspection of CCD spectra, obtained as

above on Dec. 12 UT, reveals that SN 2002jj (IAUC 8026) is a supernova of type Ic, resembling SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11) one month past maximum brightness. The spectrum of SN 2002jm (IAUC 8030) is similar to that of the subluminous type-Ia SN 1991bg (Filippenko et al. 1992, A.J. 104, 1543) at ∼ 3 weeks past maximum light, except that SN 2002jm lacks the peculiar Na I D absorption seen in SN 1991bg.

2002 December 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green